OpenStudy (anonymous):

I don't really know how to type this, but here it goes: log(base5) 4-log(base5) (x+3)= log(base5) 3?

OpenStudy (watchmath):

go ahead :)

OpenStudy (anonymous):

\[log_5(4)-log_5(x+3)=log_5(3)\]

OpenStudy (anonymous):

just wanted to type it.

OpenStudy (anonymous):

ok fine

OpenStudy (anonymous):

yeah thats it

OpenStudy (anonymous):

\[log_5(4)-log_5(3)=log_5(x+3)\] \[log_5(\frac{4}{3})=log_5(x+3)\] \[\frac{4}{3}=x+3\]

OpenStudy (anonymous):

so i guess \[x=-\frac{5}{3}\]

OpenStudy (anonymous):

you switch the 3 and the (x-3?

OpenStudy (anonymous):

only log step was using \[log(A)-log(B)=log(\frac{A}{B})\] from right to left.

OpenStudy (anonymous):

oh i just added \[log_5(x+3)\] to both sides so it would be positive. easier that way

OpenStudy (anonymous):

sorry it was adding my mistake

OpenStudy (anonymous):

added \[log_5(x+3)\] to both sides and subtracted \[log_5(3)\] from both sides so i had the variable by itself on one side.

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

clear thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

thank you for your time.